OpenStudy (anonymous):

you are making a website with a budget of $19000$ you can appoint writererw with an hourly wage of 50$/hr and designerw at 100$/hr. when xwriter work with y designers the project will be completed in 11+ 400/(xy) hrs what is the shortest time in which the project can be completed?

OpenStudy (anonymous):

I know this is a lagrange multiplier question and I have this so far: equation: F(x,y) = 11 +400/xy constraint G(x,y) = 500x + 1100y +20000y^-1+ 40000x^-1 -19000

OpenStudy (anonymous):

I'm not sure if i'm on the right track

OpenStudy (anonymous):

Fx= -400x^-2y^-1 - 500u - 40000x^-2u=0 Fx= -400x^-1y^-2 - 1100u - 20000y^-2u=0 is that right?